# Language in the Flow of Time: Time-Series-Paired Texts Weaved into a Unified Temporal Narrative

### v1

@@ 4 Texts as Time Series
Guided by the above insights, this section introduces our proposed framework, Texts as Time Series (TaTS). TaTS integrates concurrent texts by transforming them into auxiliary variables and seamlessly plugging the text-augmented time series into any existing time series model. Our approach effectively captures the temporal dynamics and interconnections between the time series and concurrent texts by unifying them within a single time series model. An overview of the proposed TaTS is illustrated in Figure 3.
@@ 4.2 Framework and Training Objectives
Our proposed TaTS framework builds on the insight that concurrent texts behave similarly to additional variables and, therefore, can be treated in a comparable manner. As shown in Figure 3.
Transforming Concurrent Texts into Variables. Given the dataset $\mathcal{D}=\{\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\},\mathbf{S}=\{s_{1},s_{2},\dots,s_{T}\}\}$, we first embed the texts using a text encoder $\mathcal{H}_{\text{text}}$ to obtain text embeddings $\mathbf{E}=\{e_{1},e_{2},\dots,e_{T}\}\in\mathbb{R}^{d_{\text{text}\times T}}$. Since the text embedding dimension $d_{\text{text}}$ is typically much larger than the number of variables in the time series, we reduce the dimensionality of the text embeddings by applying a Multi-Layer Perceptron (MLP), mapping them into a lower-dimensional space.
@@
where $\mathbf{X}_{t,i}$ and $\widehat{\mathbf{X}}_{t,i}$ represent the ground truth and predicted values of the $i^{\textrm{th}}$ variable at time step $t$, respectively.
@@ 5 Experiment
Table 2: Performance of time-series forecasting task with concurrent texts. Compared to numerical-only unimodal modeling and MM-TSFLib, our TaTS framework consistently and significantly enhances existing time series models to effectively handle time series with concurrent texts. The results are averaged across all prediction lengths. Full results are reported in Appendix C.2. Promotion denotes the percentage reduction in MSE or MAE achieved by TaTS compared to the best-performing baseline.
@@
In this section, we empirically validate the effectiveness of the proposed TaTS framework through extensive experiments conducted on multiple datasets using a variety of existing time series models. Specifically, we demonstrate that TaTS can be seamlessly integrated with existing time series models to effectively handle time series with paired texts. In particular, we use GPT2 (Radford et al., 2019) encoder to embed the paired texts. We also validate the performance of TaTS with other text encoders across different datasets in section 5.2 and Appendix C.6.
Datasets. We evaluate our framework on 9 real-world datasets (Liu et al., 2024a), spanning diverse domains such as environmental sciences, social systems, and infrastructure. The datasets have sample frequencies ranging from daily to weekly and monthly. Further details about the datasets can be found in Appendix B.1.
Time Series Models and Baselines. To demonstrate the compatibility of TaTS with existing time series models, we integrate TaTS with 9 widely used models across different categories, including (i) Transformer-based models: iTransformer (Liu et al., 2024c), PatchTST (Nie et al., 2023), Crossformer (Zhang & Yan, 2023), Autoformer (Wu et al., 2021), Informer (Zhou et al., 2021) and Transformer (Vaswani et al., 2017). (ii) Linear models: DLinear (Zeng et al., 2023). (iii) Frequency-based models: FEDformer (Zhou et al., 2022b), FiLM (Zhou et al., 2022a).
For each time series model, we compare our TaTS framework against two baselines: (i) Numerical-only uni-modal modeling, which ignores the paired texts and utilizes only the numerical time series with the given time series model. (ii) MM-TSFLib  (Liu et al., 2024a), a recently proposed library for multimodal time series forecasting that performs a linear interpolation between the output of the time series model and the text embeddings, treating the texts as a bag-of-words.
Metrics. We evaluate the performance of multimodal time series modeling using MSE, MAE, RMSE, MAPE, and MSPE. Due to space constraints, we report only the MSE and MAE results in the main paper, while the results for the other metrics are provided in Appendix C.
@@ 5.1 Main Results
Settings. Regardless of whether numerical-only models, MM-TSFLib, or TaTS are used, all models are trained to minimize the MSE loss as defined in Equation 7. Implementation details are provided in Appendix B.4.
Table 2 presents the performance results for the time series forecasting task. For each dataset and time series model, we report the average performance across four different prediction lengths, and the full results for each prediction length are provided in Appendix C. For a dataset with relatively few samples, such as Economy, we perform short-term forecasting with prediction lengths of $\{6,8,10,12\}$. In contrast, for a dataset with a larger number of samples, such as Environment, we perform long-term forecasting with prediction lengths of $\{48,96,192,336\}$. From the results, compared to uni-modal modeling using only numerical values or the state-of-the-art baseline MM-TSFLib, our TaTS consistently achieves the best performance across all datasets. Notably, by plugging in TaTS to various existing time series models, it achieves an average performance improvement of over $5\%$ on 6 out of 9 datasets and delivers a remarkable performance boost of over $30\%$ on the Environment dataset, which contains the largest number of samples among all datasets. The results also demonstrate that TaTS is highly compatible with a wide range of existing time series forecasting models, including transformer-based models, linear models, and frequency-based models, consistently delivering performance improvements across all of them in both long-term forecasting and short-term forcasting tasks. We provide a visualization of the performance boost in Appendix C.1, showcasing the improvements achieved by different time series models on each dataset.
Table 3: Performance of time-series imputation task with concurrent texts. Full results are reported in Appendix C.3. Promotion: percentage reduction in MSE or MAE achieved by TaTS compared to the best-performing baseline.
@@
We also evaluate the performance of our TaTS framework on the imputation task using the Climate, Economy, and Traffic datasets, each with an imputation length of 24. Since the original MM-TSFLib only supports forecasting tasks, we extend it by applying a similar linear interpolation approach to serve as a baseline. For this evaluation, we select one representative time series model from each category and present the results in Table 3. The results demonstrate that TaTS consistently enhances the imputation capabilities of existing time series models, achieving performance improvements of up to $30\%$ compared to baseline methods.
The above results show that leveraging the text modality provides significant benefits over numerical-only unimodal modeling when paired texts are available. Inspired by chronological textual resonance, our proposed TaTS framework effectively captures the temporal properties encoded in paired texts, achieving notable improvements over baseline methods that disregard the positional information inherent in the texts paired with time series.
Figure 4: Further analysis of our TaTS framework. (a) Learning rate sensitivity: TaTS maintains stable performance across different learning rates (full results in Appendix C.4). (b) Text Projection Dimension sensitivity: TaTS remains robust across varying $d_{\text{mapped}}$ (full results in Appendix C.5). (c) Varying text encoder: TaTS consistently outperforms baselines across different text encoders (full results in Appendix C.6). (d) Efficiency: TaTS introduces only a minor parameter increase ($\sim 1\%$) but significantly improves the performance according to Table 2.
@@ 5.2 Further Analysis
Hyperparameter Sensitivity. We perform hyperparameter studies to evaluate the impact of (i) the learning rate and (ii) $d_{\text{mapped}}$, the dimension to which high-dimensional text embeddings are projected by the MLP, as defined in Equation 4. The results are presented in Figure 4, subfigures (a) and (b), with full results available in Appendix C.4 and Appendix C.5. The findings indicate that TaTS maintains robust performance across different choices of the learning rate and the text projection dimension $d_{\text{mapped}}$.
@@
Ablation with Different Text Encoders in TaTS. While GPT-2 was used as the primary text encoder in our main experiments to demonstrate the effectiveness of TaTS, we further evaluate the performance of TaTS with different text encoders, including BERT (Devlin et al., 2019), GPT-2 (Radford et al., 2019), and LLaMA2 (Touvron et al., 2023). We utilize the official implementations available on Hugging Face and present the results in Figure 4, with full results provided in Appendix C.6. The results show that TaTS remains robust across different text encoders and consistently outperforms both the uni-modal and MM-TSFLib baselines. Notably, as the size of the language models used in TaTS increases from 110M (BERT) to 1.5B (GPT-2) and further to 7B (LLaMA2), we observe a slight improvement in performance. Investigating the relationship between the text encoder size and TaTS’s effectiveness remains an open direction for the future research.
Figure 5: Efficiency comparison on Environment dataset. While TaTS incurs a slight increase in training time due to the augmentation of auxiliary variables, it significantly enhances forecasting accuracy. Full results are in Appendix C.7.
@@
Computational Overhead vs. Performance Gain. We also evaluate the efficiency of our proposed TaTS by measuring the training time per epoch and the total number of model parameters. Figure 4 (d) presents the total number of parameters for the best-performing models in our forecasting experiments. TaTS introduces only a lightweight three-layer MLP to project high-dimensional text embeddings into a lower-dimensional space, adding a minimal number of parameters compared to the original time series models. As a result, the overall parameter count increases by only about 1%. Due to the inclusion of augmented time series with auxiliary variables from paired texts, the training time per epoch increases slightly, as shown in Figure 5, with average performance of each framework indicated by cross markers. Full results for all datasets are provided in Appendix C.7. However, this marginal efficiency trade-off ($\sim 1\%$ in terms of number of learnable parameters and $\sim 8\%$ in terms of training time) leads to significant improvements ($\sim 14\%$) in forecasting performance, demonstrating the effectiveness of TaTS in enhancing time series modeling with paired texts.
@@ 6 Related Work
Numerical-only Time Series Modeling. Recently, various deep learning models have been developed for time series analysis, which can be broadly categorized into three categories. (1) Patch-based models. PatchTST (Nie et al., 2023) segments time series into subseries-level patches to capture dependencies, while Crossformer (Zhang & Yan, 2023) employs a two-stage attention mechanism to model both cross-time and cross-variable dependencies efficiently. Autoformer (Wu et al., 2021) introduces decomposition blocks to separate seasonal and trend-cyclical components. (2) Global representation models. iTransformer (Liu et al., 2024c) utilizes attention over global series representations to capture multivariate correlations. Informer (Zhou et al., 2021) reduces self-attention complexity using ProbSparse self-attention for improved efficiency. Dlinear (Zeng et al., 2023) demonstrates that simple linear regression in the raw space can perform competitively on MTS tasks. (3) Frequency-aware models. FEDformer (Zhou et al., 2022b) represents series through randomly selected Fourier components, while FiLM (Zhou et al., 2022a) enhances representations with frequency-based layers to reduce noise and accelerate training. In this work, our proposed TaTS is compatible with all of the models listed above.
@@
Time Series with other data sources. Real-world data often coexists and coevolve with auxiliary information (Li et al., 2023; Ban et al., 2024), such as various modalities (Li et al., 2024c; Fu et al., 2022; Zhu et al., 2024; Zheng et al., 2023, 2019; Li et al., 2024b) and heterogeneous views (Zheng et al., 2024c, 2022; Li et al., 2022, 2024a). In time series analysis, effectively utilizing this auxiliary information remains an emerging and nascent area. In the financial domain, several early works have explored integrating time series with textual data, albeit not in a timestamp-aligned manner, or often leveraging general machine learning models rather than time series-specific architectures. For example, StockNet (Xu & Cohen, 2018) uses a VAE-like model for chaotic stock-text data, while (Rodrigues et al., 2019) fuses time series with a single event-related document. BoEC (Farimani et al., 2021) applies a bag-of-economic-concepts approach, and Dandelion (Zhou et al., 2020) leverages multimodal attention for feature aggregation from multiple text sources. Some works also explored time series with vision information (Liu & Cai, 2012; Gerard et al., 2023; Lütjens et al., 2024) for spatio-temporal analysis. Recently, Time-MMD (Liu et al., 2024a) constructs a dataset of time series paired with parallel text, covering multiple domains, which we used in our main experiments.
@@
Large Language Models on Time Series. The rapid advancement of Large Language Models (LLMs) and pre-trained foundation models (Jing et al., 2024b; Fu et al., 2024a; Zheng et al., 2024d) has inspired a new line of research that transforms time series into natural language representations, enabling LLMs to perform downstream tasks (Sui et al., 2024; Zhang et al., 2024a). While these approaches demonstrate strong generalization capabilities due to the power of LLMs (Gruver et al., 2023; Cao et al., 2024; Jin et al., 2024; Xue & Salim, 2024), they also inherit limitations such as hallucination (Huang et al., 2023; Zou et al., 2024) and context length constraints (Wang et al., 2024a; Liu et al., 2024b). Notably, a recent work (Tan et al., 2024) suggests that replacing complex LLM architectures with basic attention layers does not degrade the performance.

### v2

@@ 4 Texts as Time Series
This section introduces our proposed framework, Texts as Time Series (TaTS). TaTS integrates concurrent texts by transforming them into auxiliary variables and seamlessly plugging the text-augmented time series into any existing time series model. Our approach effectively captures the temporal dynamics and interconnections between the time series and concurrent texts by unifying them within a single time series model. An overview of the proposed TaTS is illustrated in Figure 3.
@@ 4.2 Framework and Training Objectives
Our proposed TaTS framework builds on the insight that concurrent texts behave similarly to additional variables and, therefore, can be treated in a comparable manner.
Transforming Concurrent Texts into Variables. Given the dataset $\mathcal{D}=\{\mathbf{X}=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\},\mathbf{S}=\{s_{1},s_{2},\dots,s_{T}\}\}$, we first embed the texts using a text encoder $\mathcal{H}_{\text{text}}$ to obtain text embeddings $\mathbf{E}=\{e_{1},e_{2},\dots,e_{T}\}\in\mathbb{R}^{d_{\text{text}\times T}}$. Since the text embedding dimension $d_{\text{text}}$ is typically much larger than the number of variables in the time series, we reduce the dimensionality of the text embeddings by applying a Multi-Layer Perceptron (MLP), mapping them into a lower-dimensional space of reduced dimensionality $d_{\text{mapped}}$.
@@
where $\mathbf{X}_{t,i}$ and $\widehat{\mathbf{X}}_{t,i}$ represent the ground truth and predicted values of the $i^{\textrm{th}}$ variable at time step $t$, respectively. TaTS for imputation follows a similar process and is omitted here due to space constraints. Detailed algorithms for both forecasting and imputation are provided in Appendix C.
To conclude this section, we note that while some existing literatures make similar but limited attempts to model covariates in multivariate time series or multiple time series. Our TaTS framework differentiates from those approaches by unifying multi-modality using deep learning architectures. In Section 5, we compare TaTS with these existing approaches.
@@ 5 Experiment
Table 2: Performance of time-series forecasting task. Compared to numerical-only unimodal modeling and MM-TSFLib, our TaTS framework consistently and significantly enhances existing time series models to effectively handle time series with paired texts. The results are averaged across all prediction lengths. Full results are reported in Appendix E.3. Promotion (positive or negative) denotes the percentage reduction (or increase) in MSE or MAE achieved by TaTS compared to the best-performing baseline.
@@
In this section, we empirically validate the effectiveness of the proposed TaTS framework through extensive experiments conducted on multiple datasets using a variety of existing time series models. In particular, we use GPT2 [43] encoder to embed the paired texts. We also validate the performance of TaTS with other text encoders across different datasets in section 5.2 and Appendix E.7. Implementation details are provided in Appendix D.4.
Datasets. We evaluate our framework on 9 real-world datasets [35], spanning diverse domains such as environmental sciences, social systems, and infrastructure. The datasets have sample frequencies ranging from daily to weekly and monthly. Further details about the datasets are in Appendix D.1.
Time Series Models and Baselines. To demonstrate the compatibility of TaTS with existing time series models, we integrate TaTS with 9 widely used models across different categories, including (i) Transformer-based models: iTransformer [37], PatchTST [39], Crossformer [67], Autoformer [56], Informer [75] and Transformer [50]. (ii) Linear models: DLinear [63]. (iii) Frequency-based models: FEDformer [77], FiLM [76]. We also compare our TaTS framework with (a) traditional covariate-based methods: N-BEATS [41], N-HiTS [4]; (b) convolution-based method TCN [1]; (c) a recent multimodal time series foundation model ChatTime [51].
For each time series model, we compare our TaTS framework against two baselines: (i) Numerical-only uni-modal modeling, which ignores the paired texts and utilizes only the numerical time series with the given time series model. (ii) MM-TSFLib  [35], a recently proposed library for multimodal time series forecasting that performs a linear interpolation between the output of the time series model and the text embeddings, treating the texts as a bag-of-words.
Metrics. We evaluate the performance of multimodal time series modeling using MSE, MAE, RMSE, MAPE, and MSPE. Due to space constraints, we report only the MSE and MAE results in the main paper, while the results for the other metrics are provided in Appendix E.
@@ 5.1 Main Results
Time Series Forecasting. Table 2 presents the performance results for the time series forecasting task. For each dataset and time series model, we report the average performance across four different prediction lengths, and the full results for each prediction length are provided in Appendix E. For datasets with relatively few samples, such as Economy, we perform short-term forecasting with prediction lengths of $\{6,8,10,12\}$. In contrast, for datasets with a larger number of samples, such as Environment, we perform long-term forecasting with prediction lengths of $\{48,96,192,336\}$. From the results, compared to uni-modal modeling or the state-of-the-art baseline MM-TSFLib, our TaTS consistently achieves the best performance across all datasets. Notably, by plugging in TaTS to various existing time series models, it achieves an average performance improvement of over $5\%$ on 6 out of 9 datasets and delivers a remarkable performance boost of over $30\%$ on the Environment dataset, which contains the largest number of samples among all datasets. The results also demonstrate that TaTS is highly compatible with a wide range of existing time series forecasting models, including transformer-based models, linear models, and frequency-based models, consistently delivering performance improvements across all of them in both long-term forecasting and short-term forecasting tasks. We provide a visualization of the performance boost in Appendix E.2, showcasing the improvements achieved by different time series models on each dataset.
Table 3: Performance of time-series imputation task with concurrent texts. Full results are reported in Appendix E.4. Promotion: percentage reduction in MSE or MAE achieved by TaTS compared to the best-performing baseline.
@@
Time Series Imputation. We also evaluate the performance of our TaTS framework on the imputation task using the Climate, Economy, and Traffic datasets, each with an imputation length of 24. Since the original MM-TSFLib only supports forecasting tasks, we extend it by applying a similar linear interpolation approach to serve as a baseline. For this evaluation, we select one representative time series model from each category and present the results in Table 3. The results demonstrate that TaTS consistently enhances the imputation capabilities of existing time series models, achieving improvements of up to $30\%$ compared to baseline methods.
From the above results, it is evident that leveraging the text modality provides significant benefits over numerical-only unimodal modeling when paired texts are available. Our proposed TaTS framework effectively captures the temporal properties encoded in paired texts, achieving notable improvements over baseline methods that disregard the positional information inherent in the texts paired with time series.
Table 4: TT-Wasserstein measure of Time-MMD [35] datasets and their random shuffle.
Dataset Monthly Sampled Weekly Sampled Daily Sampled
Agriculture Climate Economy Security Social Good Traffic Energy Health Environment
Ratio of Original to Shuffled TT-Wasserstein (%) 26.8 72.4 22.3 91.5 38.2 34.6 97.1 85.6 83.6
Avg. Improvement of TaTS over uni-modal (%) 20.70 18.03 64.80 4.05 10.99 16.78 3.62 19.51 36.00
TaTS Benefits from Better Alignment (i.e., lower TT-Wasserstein). We compute the average forecasting improvement of TaTS compared to numerical-only modeling, as well as the ratio of TT-Wasserstein between original and shuffled Time-MMD datasets. The results, shown in Table 4, reveal that, within the same sampling frequency, lower TT-Wasserstein original-shuffled ratios tend to correlate with larger performance gains from TaTS (except for Climate). In other words, for paired time series and texts that have stronger CTR, i.e., more relevant, our TaTS could improve performance more. Thus, TT-Wasserstein can indicate the usefulness of paired texts and the potential effectiveness of TaTS when text quality and cross-modal alignment are uncertain.
Figure 4: Further analysis of our TaTS framework. (a) Learning rate sensitivity: TaTS maintains stable performance across different learning rates (full results in Appendix E.5). (b) Text Projection Dimension sensitivity: TaTS remains robust across varying $d_{\text{mapped}}$ (full results in Appendix E.6). (c) Varying text encoder: TaTS consistently outperforms baselines across different text encoders (full results in Appendix E.7). (d) Efficiency: TaTS introduces only a minor parameter increase ($\sim 1\%$) but significantly improves the performance according to Table 2.
Table 5: TaTS compared several baselines.
Dataset Climate Security Traffic
Methods MSE MAE MSE MAE MSE MAE
iTransformer [2024c] 1.127 0.843 113.57 5.698 0.203 0.228
N-BEATS [2020] 1.123 0.894 131.91 6.866 0.352 0.409
N-HiTS [2022] 1.114 0.881 135.41 6.896 0.295 0.372
TCN [2018] 1.137 0.875 129.61 6.801 0.779 0.775
ChatTime [2025] 1.507 1.007 130.75 6.854 0.369 0.435
iTransformer + TaTS (ours) 1.020 0.797 107.11 4.856 0.174 0.218
Comparison with Various Baselines. Besides numerical-only time series models and the MM-TSFLib framework, we compare our TaTS with covariate-based methods N-BEATS [41] and N-HiTS [4], convolution-based TCN [1], and the recent multimodal time series foundation model ChatTime [51]. Although these methods are not specifically designed for time series with paired texts, we adapt them by feeding both time series and text embeddings to covariate-based methods and TCN, while for ChatTime, we concatenate all paired texts into a single long text and perform zero-shot inference. The results are shown in Table 5. The results show that covariate-based and convolution-based models perform comparably or worse than iTransformer, and consistently underperform compared to our TaTS. Although ChatTime achieves competitive results even in a zero-shot setting, highlighting the value of integrating texts when high-quality texts are available, our TaTS outperforms it through supervised learning.
@@ 5.2 Further Analysis
Hyperparameter Sensitivity. We perform hyperparameter studies to evaluate the impact of (i) the learning rate and (ii) $d_{\text{mapped}}$, the dimension to which high-dimensional text embeddings are projected by the MLP, as defined in Equation 5. The results are presented in Figure 4, subfigures (a) and (b), with full results available in Appendix E.5 and Appendix E.6. The findings indicate that TaTS maintains robust performance across different choices of the learning rate and the text projection dimension $d_{\text{mapped}}$.
@@
Ablation with Different Text Encoders in TaTS. While GPT-2 was used as the primary text encoder in our main experiments to demonstrate the effectiveness of TaTS, we further evaluate the performance of TaTS with different text encoders, including BERT [7], GPT-2 [43], and LLaMA2 [49]. We utilize the official implementations available on Hugging Face and present the results in Figure 4, with full results provided in Appendix E.7. The results show that TaTS remains robust across different text encoders and consistently outperforms both the uni-modal and MM-TSFLib baselines. Notably, as the size of the language models used in TaTS increases from 110M (BERT) to 1.5B (GPT-2) and further to 7B (LLaMA2), we observe a slight improvement in performance. Investigating the relationship between the text encoder size and TaTS’s effectiveness remains an open direction for future research.
Table 6: Ablation study of TaTS on text-shuffled datasets.
Dataset Climate Security Traffic
Methods MSE MAE MSE MAE MSE MAE
iTransformer + TaTS + original data 1.020 0.797 107.11 4.856 0.174 0.218
iTransformer + TaTS + corrupted data 1.147 0.831 116.48 5.714 0.208 0.270
iTransformer + uni-modal + original data 1.127 0.843 113.57 5.698 0.203 0.228
Ablation on Corrupted Datasets. To validate that the performance gains stem from useful textual cues, we conducted corruption experiments where the textual information was randomly shuffled across timestamps. All other settings are unchanged, and the results are in Table 6. The results demonstrate that when textual alignment is destroyed, the model’s performance drops significantly to matching or even being worse than the uni-modal baseline. This is expected, as randomly shuffled text acts as noise rather than a meaningful signal.
@@
Computational Overhead vs. Performance Gain. We also evaluate the efficiency of our proposed TaTS by measuring the training time per epoch and the total number of model parameters. Figure 4 (d) presents the total number of parameters for the best-performing models in our forecasting experiments. TaTS introduces only a lightweight three-layer MLP to project high-dimensional text embeddings into a lower-dimensional space, adding a minimal number of parameters compared to the original time series models. As a result, the overall parameter count increases by only about 1%.
Due to the inclusion of augmented time series with auxiliary variables from paired texts, the training time per epoch increases slightly, as shown in Figure 5, with average performance of each framework indicated by cross markers. Full results for all datasets are provided in Appendix E.8. However, this marginal efficiency trade-off ($\sim 1\%$ in terms of number of learnable parameters and $\sim 8\%$ in terms of training time) leads to significant improvements ($\sim 14\%$) in forecasting performance, demonstrating the effectiveness of TaTS in enhancing time series modeling with paired texts.
Figure 5: Efficiency comparison on Environment dataset. While TaTS incurs a slight increase in training time due to the augmentation of auxiliary variables, it significantly enhances forecasting accuracy. Full results in Appendix E.8.
@@ 6 Related Work
Numerical-only Time Series Modeling. Recently, various deep learning models have been developed for time series analysis, which can be broadly categorized into three categories. (1) Patch-based models. PatchTST [39] segments time series into subseries-level patches to capture dependencies, while Crossformer [67] employs a two-stage attention mechanism to model both cross-time and cross-variable dependencies efficiently. Autoformer [56] introduces decomposition blocks to separate seasonal and trend-cyclical components. (2) Global representation models. iTransformer [37] utilizes attention over global series representations to capture multivariate correlations. Informer [75] reduces self-attention complexity using ProbSparse self-attention for improved efficiency. Dlinear [63] demonstrates that simple linear regression in the raw space can perform competitively on MTS tasks. (3) Frequency-aware models. FEDformer [77] represents series through randomly selected Fourier components, while FiLM [76] enhances representations with frequency-based layers to reduce noise and accelerate training. In this work, our proposed TaTS is compatible with all of the models listed above.
@@
Time Series with other data sources. Real-world data often coexists and coevolve with auxiliary information [29, 2], such as various modalities [32, 9, 78, 70, 68, 31] and heterogeneous views [73, 69, 28, 30]. In time series analysis, effectively utilizing this auxiliary information remains an emerging and nascent area. In the financial domain, several early works have explored integrating time series with textual data, albeit not in a timestamp-aligned manner, or often leveraging general machine learning models rather than time series-specific architectures. For example, StockNet [60] uses a VAE-like model for chaotic stock-text data, while [45] fuses time series with a single event-related document. BoEC [8] applies a bag-of-economic-concepts approach, and Dandelion [74] leverages multimodal attention for feature aggregation from multiple text sources. Some works also explored time series with vision information [34, 11, 38] for spatio-temporal analysis. Recently, Time-MMD [35] constructs a dataset of time series paired with parallel text, covering multiple domains, which we used in our main experiments.
@@
Large Language Models on Time Series. The rapid advancement of Large Language Models (LLMs) has inspired a new line of research that transforms time series into natural language representations, enabling LLMs to perform downstream tasks [66]. While these approaches demonstrate strong generalization capabilities due to the power of LLMs [12, 3, 16, 61], they also inherit limitations such as hallucination [14] and context length constraints [52, 36]. Notably, a recent work [48] suggests that replacing complex LLM architectures with basic attention layers does not necessarily degrade the performance.
In Appendix B, we also discuss available preprints that are related to this research.